OpenStudy (anonymous):

MEDAL! Find all the zeros of the equation: x^4-6x^2-7x-6=0

Miracrown (miracrown):

I can't see any way to factor it ... Maybe it can be graphed?

OpenStudy (jjuden):

\[x=\frac{ -\frac{ 1 }{ 2 }+\frac{ i }{ 2 }\pm \frac{ \sqrt{3}(1+1) }{ 2 } }{ 1 }-i\]

Miracrown (miracrown):

If there is a "nice" root, we can use the factors of the last term +-6, +-3, +-2, +-1 divided by the factors of the first term (+-1) I'd suggest using synthetic division since we have x^4 it can't be very large

Miracrown (miracrown):

It's easy to plug in +-1 to test 1: 1 - 6 -7 - 6 = -18 -1: 1 -6 +7 -6 = -4 What do we get if you plug +-2 in ?

OpenStudy (anonymous):

I am so confused right now.... Where do I plug that in?

Miracrown (miracrown):

synthetic division with +-2 ?

Miracrown (miracrown):

is there a remainder ?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

x^3-2x^2-2x-3

Miracrown (miracrown):